# Towards Understanding the Benefits of Neural Network Parameterizations in Geophysical Inversions: A Study With Neural Fields

### v1

@@ IV-B Apply Positional Encoding to Recovery the Higher-frequency Variances
Figure 9: a) is the true model for Case 2. b) is the proposed NFs-Inv inversion result with an identity $\gamma$. c) is the proposed NFs-Inv inversion result with a Gaussian positional encoding where $B\in R^{128\times 2}$ in eqn. 7.
Figure 10: a) is the true model for Case 4. b) is the proposed NFs-Inv result with a linear positional encoding where m = 8 in eqn. 6. c) is the proposed NFs-Inv result with an identity $\gamma$.
As mentioned before, mapping the inputs to a higher dimensional space using high-frequency functions before passing them to the network enables better fitting of high-frequency variation data. That is why we observe a better recovery of the heterogeneous object in the seismic tomographic example. The fact that we observe the deteriorated recovery of the model in Case 4 reflects that the choice of positional encoding should align with the ability of the geophysical survey to resolve the subsurface model. The resolution of a survey depends upon the underlying physics of the problem, survey geometry, field complexity, etc. For example, a DCR survey typically has less resolution than a borehole seismic survey. A more practical guide for choosing positional encodings is an area for future research and will be informed by conducting more trials on different subsurface models and survey methods.
@@ V Analysis using the singular value decomposition analysis of the Jacobian of the Neural Network
While deep learning has demonstrated remarkable success across a growing range of tasks, a comprehensive understanding of the underlying principles driving these achievements remains limited. Many empirical-based methods (“scientific method”) have been developed for studying the underlying mechanism [36], [37], [26]. For example, Kadkhodaie et al. [26] find that deep denoisers have inductive biases toward learning a geometry adaptive harmonic basis, and it could potentially lead to solving the generalization puzzle in many large ML models. The authors acknowledge that there must be priors from NN structures and the optimization methods, and the authors refer to it as inductive bias. For clarification, we use implicit bias instead of inductive bias in our context to avoid potential ambiguity since there are some papers that refer to inductive bias in CNN specifically as translational invariance [53]. Basically, Kadkhodaie et al. [26] state that since a DNN can be considered as a locally linear function, conducting an eigendecomposition analysis (SVD analysis) of the Jacobian of the neural network concerning the input can reveal what has been learned by the generative models. Namely, since these left-singular vectors form the basis of the output, the features of the basis can indicate what has been imposed by the implicit bias during the training process. They find that the learned bases have oscillating harmonic structures along contours and in homogeneous regions. They name these bases as geometry-adaptive harmonic bases (GAHBs).
In our work, we also find that the Jacobian has a geometry-adaptive harmonic basis after inversion. Note that Kadkhodaie et al. [26] conducted an SVD analysis of the Jacobian of the output w.r.t. the input. In their supervised learning setting, the inputs of the model are different for different input images. However, in our case, since the inputs are always the fixed values of the coordinates of each cell, we conduct an SVD analysis of the Jacobian of the output w.r.t. the weights of the NN $f$. Since the choice of the positional encoding may introduce an additional bias (See Section IV-B), we do not employ positional encoding for the models presented in this subsection (i.e., $\gamma$ is an identity function.) to avoid potential ambiguity. The SVD of the Jacobian is
@@
The left-singular vectors (the columns of $U$) have the dimensionality of the geophysical model and the right-singular vectors (the columns of $V$) have the dimensionality of the number of weights in the neural network. $\Lambda$ is a diagonal matrix whose elements are the singular values $\lambda_{i}$. We focus our analysis on the left-singular vectors as these can be thought of as the learned basis for our geophysical model
@@
From the plots of the eigenvalues for all 4 examples, we find that the eigenvalues are sparse. Namely, the first few eigenvalues are significantly larger than the remaining eigenvalues. In Case 1 and Case 3, which are both relatively simple, there is a reduction of nearly a factor of 10 between the first and tenth singular values. For Case 2, which is more complex, there is a reduction of a factor of around 3 between the first and tenth singular values. This observation agrees with the finding in the analysis of Jacobian in the diffusion models [26] where they also find this shrinkage of the eigenvalues.
These examples are evidence that there exists an implicit bias, which has been illustrated in supervised learning models and may also be utilized in the test-time learning models as well. Moreover, this analysis can partly explain the reason why we can observe a better recovery of the subsurface physical properties when we re-parameterize the subsurface model using the weights of a DNN. By forming these adaptive bases, we have the chance to overcome the problems of predicting unwanted artifacts due to sensitivity, such as near-electrode artifacts or the concentration along the ray path.
@@ VI Discussion and Future Work
Further work, such as the applications in the field data and other theoretical analyses of the implicit bias is still in progress. It has been shown that using the adaptive grids can improve the performance of the neural fields methods [54], [8], so testing the adaptive grid scheme for geophysical inversions is a possible future research direction.
The run-time needed for convergence of the NFs-Inv algorithm heavily depends on the time required to run the forward simulation. This is a limitation of the current implementation of DC resistivity forward modelling, which runs on CPUs. As a result, improving the efficiency of the forward simulation would decrease the time cost of the proposed method.
The results shown in Section III are state-of-the-art which we established by testing the performance of different positional encoding functions. We found that in the cases where the subsurface model is relatively simple such as Cases 1 &\&& 3, adding a positional encoding function where $h$ is large will deteriorate the inversion results. However, the results of Case 2 show that when the model is more complex and we use a geophysical method with sufficient sensitivity, adding positional encoding can help us to recover sharper boundaries and finer details in the model. There is still some work left on exploring the effect of the choice of $\gamma$ function on the proposed method.
@@ Acknowledgments
This research was supported by funding from the NSERC Discovery Grants Program and the Mitacs Accelerate Program. We would like to thank UBC GIF group members for their invaluable guidance and advice on the usage of SimPEG.

### v2

@@ IV-B Apply Positional Encoding to Recover the Higher-frequency Variation
Figure 9: (a) is the true model for Case 2. (b) is the proposed NFs-Inv inversion result with an identity $\gamma$. (c) is the proposed NFs-Inv inversion result with a Gaussian positional encoding where $B\in R^{128\times 2}$ in eqn. 7.
Figure 10: ((a) is the true model for Case 4. (b) is the proposed NFs-Inv result with a linear positional encoding where m = 8 in eqn. 6. (c) is the proposed NFs-Inv result with an identity $\gamma$.
As mentioned before, mapping the inputs to a higher-dimensional space using high-frequency functions prior to passing them to the network enables better fitting of high-frequency variation data. That is why we observe a better recovery of the heterogeneous object in the seismic tomographic example. The fact that we observe the deteriorated recovery of the model in Case 4 illustrates that the choice of positional encoding should align with the ability of the geophysical survey to resolve the subsurface model. Namely, the positional encoding introduces an additional bias to the geophysical inverse problems, which may deteriorate the results in some cases. The resolution of a survey depends upon the underlying physics of the problem, survey geometry, geologic complexity, etc. In the case of cross-hole seismic tomography inversion, the sensors are located on both sides of the survey domain, so there are enough data to constrain the possible solutions. The inversion results with or without the positional encoding can recover similar contours of the subsurface model. Therefore, adding a positional encoding can be helpful for recovering the high-frequency component in the subsurface model in the seismic case. In contrast, the DCR survey for Case 3 &\&& 4 only has sensors on the ground, and the survey domain is larger than Case 1 &\&& 2. As a result, Case 3 &\&& 4 is much more ill-posed than Cases 1 &\&& 2. Namely, the solution is less constrained, and there are more geologically plausible solutions that can fit the observations to the same degree. In the DCR examples, the additional bias provided by the positional encoding may cause the inversion result to have unrealistically high-frequency signals. As a result, whether we choose to employ a positional encoding should depend on the type of survey method. In terms of selecting the optimal positional encoding, we found that the Gaussian positional encoding with a small standard deviation for generating the matrix $B$ (eqn. (7)) can produce similar results as the basic positional encoding. The selection of the hyperparameters of the positional encoding depends on the complexity of the subsurface model. A more practical guide for choosing positional encodings is an area for future research and will be informed by conducting more trials on different subsurface models and survey methods.
@@ V Analysis using the singular value decomposition analysis of the Jacobian of the Neural Network
While deep learning has demonstrated remarkable success across a growing range of tasks, a comprehensive understanding of the underlying principles driving these achievements remains limited. Many empirical-based methods (“scientific method”) have been developed for studying the underlying mechanism [44], [45], [35]. For example, Kadkhodaie et al. [35] find that deep denoisers have inductive biases toward learning a geometry adaptive harmonic basis, and it could potentially lead to solving the generalization puzzle in many large ML models. The authors acknowledge that there must be priors from NN structures and the optimization methods, and the authors refer to it as inductive bias. For clarification, we use implicit bias instead of inductive bias in our context to avoid potential ambiguity since some papers refer to inductive bias in CNN specifically as translational invariance [61]. Kadkhodaie et al. [35] state that since a DNN can be considered as a locally linear function, conducting an eigendecomposition analysis (SVD analysis) of the Jacobian of the neural network concerning the input can reveal what has been learned by the generative models. Namely, since these left-singular vectors form the basis of the output, the features of the basis can indicate what has been imposed by the implicit bias during the training process. They find that the learned bases have oscillating harmonic structures along contours and in homogeneous regions. They name these bases as geometry-adaptive harmonic bases (GAHBs).
In our work, we also find that the Jacobian has a geometry-adaptive harmonic basis after inversion. Note that Kadkhodaie et al. [35] conducted an SVD analysis of the Jacobian of the output with respect to the input. In their supervised learning setting, the inputs of the model are different for different input images. However, in our case, since the inputs are always the fixed values of the coordinates of each cell, we conduct an SVD analysis of the Jacobian of the output with respect to the weights of the NN $f$. Since the choice of the positional encoding may introduce an additional bias (See Section IV-B), we do not employ positional encoding for the models presented in this subsection (i.e., $\gamma$ is an identity function) to avoid potential ambiguity. The SVD of the Jacobian is
@@
The left-singular vectors (the columns of $U$) have the dimensionality of the geophysical model, and the right-singular vectors (the columns of $V$) have the dimensionality of the number of weights in the neural network. $\Lambda$ is a diagonal matrix whose elements are the singular values $\lambda_{i}$. We focus our analysis on the left-singular vectors, as these can be thought of as the learned basis for our geophysical model.
@@
From the plots of the eigenvalues for all 4 examples, we find that the eigenvalues are sparse. Namely, the first few eigenvalues are significantly larger than the remaining eigenvalues. In Case 1 and Case 3, which are both relatively simple, there is a reduction of nearly a factor of 10 between the first and tenth singular values. For Case 2, which is more complex, there is a reduction of a factor of around 3 between the first and tenth singular values. This observation agrees with the finding in the analysis of Jacobian in the diffusion models [35], where they also find this shrinkage of the eigenvalues.
These examples are evidence that there exists an implicit bias, which has been illustrated in supervised learning models and may also be utilized in test-time learning models as well. Moreover, this analysis can partly explain the reason why we can observe a better recovery of the subsurface physical properties when we re-parameterize the subsurface model using the weights of a DNN. By forming these adaptive bases, we have the chance to overcome the problems of predicting unwanted artifacts due to sensitivity, such as near-electrode artifacts or the concentration along the ray path.
@@ VI Discussion and Future Work
Further work, such as the applications to the field data and other theoretical analyses of the implicit bias, is still in progress. It has been shown that using the adaptive grids can improve the performance of the neural fields methods [62], [11], so testing the adaptive grid scheme for geophysical inversions is a possible future research direction.
[…]
The results shown in Section III are state-of-the-art, which we established by testing the performance of different positional encoding functions. We found that in the cases where the subsurface model is relatively simple, such as Cases 1 &\&& 3, adding a positional encoding function where $h$ is large will deteriorate the inversion results. However, the results of Case 2 show that when the model is more complex and we use a geophysical method with sufficient sensitivity, adding positional encoding can help us to recover sharper boundaries and finer details in the model. There is still some work left on exploring the effect of the choice of the $\gamma$ function on the proposed method.
@@ Acknowledgments
This research was supported by funding from the NSERC Discovery Grants Program and the Mitacs Accelerate Program. We would like to thank Zahra Kadkhodaie for her elaboration of her works. We also want to thank UBC GIF group members for their invaluable guidance and advice on the usage of SimPEG. We are grateful for the careful review and constructive feedback provided by the reviewers.